f(x) = 5^x
Please help!!! how would I show on a table of coordinates and what is the the additional points?

Answer by stanbon(75887)

f(x) = 5^x
If x=-2, y=5^-2=1/25 giving you point(-2,1/25)
If x=-1, y=5^-1 =1/5 giving you point (-1,1/5)
If x=0, y=5^0=1 giving you point (0,1)
If x=1, y=5^1=5 giving you point (1,5)
If x=2, y=5^2=25 giving you point (2,25)
etc.
===============
Cheers,
Stan H